# Taming generative video models for zero-shot optical flow extraction

### v1

@@ Abstract
Extracting optical flow from videos remains a core computer vision problem. Motivated by the recent success of large general-purpose models, we ask whether frozen self-supervised video models trained only to predict future frames can be prompted, without fine-tuning, to output flow. Prior attempts to read out depth or illumination from video generators required fine-tuning; that strategy is ill-suited for flow, where labeled data is scarce and synthetic datasets suffer from a sim-to-real gap. Inspired by the Counterfactual World Model (CWM) paradigm, which can obtain point-wise correspondences by injecting a small tracer perturbation into a next-frame predictor and tracking its propagation, we extend this idea to generative video models for zero-shot flow extraction. We explore several popular architectures and find that successful zero-shot flow extraction in this manner is aided by three model properties: (1) distributional prediction of future frames (avoiding blurry or noisy outputs); (2) factorized latents that treat each spatio-temporal patch independently; and (3) random-access decoding that can condition on any subset of future pixels. These properties are uniquely present in the recently introduced Local Random Access Sequence (LRAS) architecture. Building on LRAS, we propose KL-tracing: a novel test-time inference procedure that injects a localized perturbation into the first frame, rolls out the model one step, and computes the Kullback–Leibler divergence between perturbed and unperturbed predictive distributions. Without any flow-specific fine-tuning, our method outperforms state-of-the-art models on real-world TAP-Vid DAVIS dataset (16.6% relative improvement for endpoint error) and synthetic TAP-Vid Kubric (4.7% relative improvement), despite being trained on real-world videos. Our results indicate that counterfactual prompting of controllable generative video models is a scalable and effective alternative to supervised or photometric-loss approaches for high-quality optical flow.  [^1] Project website at: https://neuroailab.github.io/projects/kl_tracing/
@@
Figure 1: We introduce a zero-shot test-time inference procedure called KL-tracing, which extracts robust optical flow and point tracking from a generative world model on challenging in-the-wild videos. In every column, the green line links the query location in the first frame (top) to the position predicted by our method in the second frame (bottom). All clips are real-world internet videos and contain phenomena that classical, appearance-based optical flow methods find challenging: (A) Newton’s cradle, where both frames have four balls in the middle, but the balls are different; the example involves physical reasoning. (B) Globe has challenging in-place object rotation and the query point is in the textureless ocean. (C) Dog weaving through occluding poles with large, rapid motion, including depth changes and motion blur. (D) Soccer tackle with fast, diagonal motion with motion blur and partial occlusion. (E) Windmill rotation where the repetitive blades and uniform sky make local matching challenging. These examples highlight the benefits of leveraging a powerful world model to extract optical flow for challenging real-world scene dynamics.
@@ 1 Introduction
Extracting motion information (optical flow) from videos is a fundamental yet open challenge in computer vision with many applications. Due to the intractable cost of obtaining ground-truth labels from real-world videos, most supervised baselines [37, 39] are trained on synthetic datasets, e.g., FlyingChairs [12], FlyingThings [27] and Sintel [6]. While invaluable for progress, these datasets cover a narrow slice of motion statistics–predominantly rigid objects, limited lighting variation, and short temporal horizons. As a result, they under-represent the long-tail of real-world phenomena such as non-rigid deformation, atmospheric effects, rapid camera shake and textureless regions. Self-supervised models [21, 36] that can be trained on real videos attempt to bridge this gap, but they often rely on task-specific heuristics, such as photometric consistency or smoothness which fail under complex lighting, occlusion, or long-range motion. Consequently, both supervised and self-supervised optical flow baselines struggle to generalize to challenging in-the-wild videos.
@@
Inspired by successes across vision and language where large general-purpose models outperform smaller task-specific ones, we explore large-scale video models as a possible solution. Trained on massive repositories of real-world data, modern video models already demonstrate strong scene understanding [5, 29, 1, 26], suggesting an implicit grasp of optical flow [3, 38]. However, prior work extracting visual intermediates such as depth and illumination from these models still required supervised fine-tuning [33, 11]. Extending that recipe to optical flow would again depend on synthetic labels, facing the same sim-to-real domain gap. This motivates our search for a zero-shot procedure that can extract accurate optical flow from off-the-shelf video models without any additional training.
@@
In fact, such a procedure has been proposed in the Counterfactual World Model (CWM) framework [3]. Zero-shot optical flow is obtained by adding a small “tracer” perturbation to the source frame and tracking how a pretrained next-frame predictor propagates the tracer to the target frame (Section 2). In practice, however, because deterministic video models like CWM can only predict a single future state, it encourages predictions that average over future possibilities, yielding perceptually blurry frames that wash out the injected tracer, leading to less precise motion estimates (Section 4.1).
@@
To overcome this, we implement the perturb-and-track method with generative video models, which generate crisp predictions as they sample from a distribution instead of regressing to a mean. This is insufficient, however, as each state-of-the-art generative video model faces its own flow extraction challenges. Stable Video Diffusion (SVD) [5] produces photorealistic frames, but its conditioning is weak and non-localized: generation is guided by a single global latent inverted from the target frame, so pixel-level edits introduce noisy differences, corrupting the extracted flow (Section 4.2). The recent Cosmos model [29] can condition generation on ground-truth patches from the target frame, but its autoregressive rollout is strictly in raster order, so these provided patches must all lie in the raster top-left region of the image, far less informative than the same number of patches distributed across the image, thus failing to accurately reconstruct the target frame (Section 4.3).
@@
Through a systematic analysis of state-of-the-art video models, we find three key properties for accurate zero-shot optical flow extraction: (1) distributional prediction of future frames (avoiding blurry or noisy outputs); (2) local tokenization, which encodes each spatio-temporal patch independently; and (3) random access decoding that can condition on any subset of future pixels. These are present in the recent Local Random Access Sequence (LRAS) architecture [24], whose patch-level conditioning and random-order decoding provide much more fine-grained controllability than previous generative video models (Section 4.4). In addition, we propose a new method, KL-tracing, that further improves upon the probe-and-track method by using the autoregressive nature of LRAS which exposes the probability distributions for predictions. KL-tracing computes the perturbation difference in the prediction logit space, thereby bypassing noisy RGB differences resulting from sampling randomness. We achieve state-of-the-art results on the challenging real-world TAP-Vid DAVIS dataset and the synthetic TAP-Vid Kubric dataset relative to supervised and unsupervised flow baselines (Section 6).
@@ 2 Related Work
Optical Flow refers to estimating the per-pixel 2D motion between a pair of video frames. Most models [37, 39] are supervised on synthetic video datasets [12, 27, 6], posing a sim-to-real gap, i.e., predominantly rigid objects, limited lighting variation, and short temporal horizons, failing to capture the long tail of real-world phenomena such as non-rigid deformation, atmospheric effects, fast camera shakes, and fine-scale textureless regions. Alternatively, self-supervised models [22, 25, 36] train on real-world videos, but often rely on task-specific architectures and heuristics such as photometric consistency and smoothness loss [22, 25, 36]. Regardless, these optical flow methods are designed for short frame gaps and relatively non-complex scenes with little motion dynamics.
@@
Point Tracking follows a set of points across longer time horizons. Most methods are supervised [9, 17, 23], often on synthetic datasets, facing the sim-to-real gap. Alternatively, self-supervised methods rely on task-specific heuristics such as smoothness losses and cycle consistency [20, 4, 34], i.e., tracking a point forward in time, then backward, should return to its original position. In contrast, we use a zero-shot method to extract flow from general-purpose self-supervised video models trained on diverse, large-scale datasets, as they better learn challenging real-world dynamics (Section 6).
@@
Deterministic video models predict a single future frame/latent  [15, 14, 16, 3, 2]. Trained with either latent or pixel $\ell_{1}/\ell_{2}$ reconstruction losses, they are implicitly optimised to output the expectation over plausible futures. Whenever the conditioning signal—a single frame $F_{1}$, multiple frames $(F_{-N}\!\ldots\!F_{1})$, or a partially masked target $F_{2}^{\text{masked}}$—can lead to multiple future possibilities, the model averages them, producing spatially blurred predictions which hurt flow estimation.
@@
Generative Video Models. Generative models can sample different image or video generations from a probabilistic distribution. Diffusion models iteratively denoise a single noise sample into a generation [18, 35, 32, 19, 5]. Autoregressive models [30, 7, 31, 8] generate outputs sequentially, modeling the joint distribution over pixels or tokens, each element conditioned on previously generated elements. However, many of these models, e.g., Cosmos [29], predict tokens in raster order and use a global tokenizer where each image patch is encoded with information from other patches. Hence, they lack fine-grained controllability which makes them challenging to use for flow extraction (Section 4.3) In contrast, a recent class of models, Local Random Access Sequence (LRAS), has several key properties which make them suitable for optical flow extraction (Section 4.4). Applying our novel test-time inference procedure to LRAS achieves state-of-the-art optical flow (Section 6).
@@ 3 Test-time inference procedure and evaluation setup for optical flow
Figure 2: Test-time inference procedure for extracting flow from a pre-trained, frozen, generative video model, based on the Counterfactual World Model (CWM) paper [3]. This involves three steps: (1) Perturbation: add a small, white-colored 2D Gaussian dot perturbation to frame 1 at the location of the point we wish to track. (2) Generate model predictions conditioned on the two frames. For CWM, Cosmos, and LRAS, we provide frame 1 and masked patches of frame 2 (Sections 4.1, 4.3, 4.4). For Stable Video Diffusion, we provide the noised latents of both frames (Section 4.2). (3) Estimate optical flow by computing the RGB difference between the clean and perturbed predictions.
We use a zero-shot procedure to extract optical flow from various pre-trained video models. For all models, we utilize the generic method template illustrated in Figure 2, based on the Counterfactual World Model (CWM) paper [3]. Step 1. Inject a small perturbation. We duplicate the initial frame $F_{1}$ and perturb it with a small “white bump" to form $\tilde{F}_{1}$, i.e., a Gaussian centered at the query location $x_{q}$ with amplitudes 255 on each RGB channel and standard deviation $\sigma$ equal 2.0. Step 2. Run model twice. Both clean and perturbed initial frames are separately forwarded through the frozen model, each time with a sparse mask of the second frame $({F}_{2}^{\text{masked}})$. Hence, two forward passes: $({F}_{1},{F}_{2}^{\text{masked}}\rightarrow{F}_{2}^{\text{pred}})$ and $(\tilde{F}_{1},{F}_{2}^{\text{masked}}\rightarrow\tilde{F}_{2}^{\text{pred}})$. Step 3. Estimate optical flow. Compute the RGB difference between the two predictions, $F_{2}^{\text{pred}}$ and $\tilde{F}_{2}^{\text{pred}}$. Then, take the arg-max to identify where the perturbation was carried to.
@@
Evaluation setup. We use TAP-Vid DAVIS [9] and Kubric [13] for evaluation. TAP-Vid DAVIS contains real-world videos with human-annotated flow and occlusion labels, while Kubric is a synthetic dataset with ground-truth labels. Metrics. We use the following metrics. (1) Average Distance (AD) (or endpoint error (EPE)), the Euclidean distance between the predicted and ground-truth flow, as well as metrics from TAP-Vid [9]: (2) Average Jaccard (AJ), the “true positives" divided by “true positives" plus “false positives" plus “false negatives", averaged over various thresholds; (3) $<\!\delta^{x}_{\text{avg}}\!$, the fraction of points that are within a threshold of the ground truth location, averaged over various thresholds, and (4) Occlusion Accuracy (OA) for predicting occluded points.
@@ 4 Model evaluations
We conduct several studies extracting optical flow from various video model classes, highlighting issues we found with each. We investigate deterministic models (Section 4.1), diffusion models (Section 4.2), and autoregressive models (Section 4.3). Finally, we identify the key properties of generative models that enable precise optical flow, and a model class that satisfies them (Section 4.4).
4.1 Study 1: Deterministic models produce blurry-averaged predictions.
Method: Deterministic video models such as the Counterfactual World Model (CWM) [3, 38] predict a single future state. CWM is a visual foundation model that can be zero-shot prompted to perform many tasks, such as predicting flow, keypoints, object segments, counterfactuals, and depth. CWM was trained with an $\ell_{2}$ loss to perform masked next-frame prediction: $({F}_{1},F_{2}^{\text{masked}}\rightarrow{F}_{2}^{\text{pred}})$. It deterministically predicts a single future state ${F}_{2}^{\text{pred}}$, as it lacks a mechanism for sampling. To extract flow from CWM, we use the procedure in Section 2 with a small, red perturbation, following [3].

### v2

@@ Abstract
Extracting optical flow from videos remains a core computer vision problem. Motivated by the recent success of large general-purpose models, we ask whether frozen self-supervised video models trained only to predict future frames can be prompted, without fine-tuning, to output flow. Prior attempts to read out depth or illumination from video generators required fine-tuning; that strategy is ill-suited for flow, where labeled data is scarce and synthetic datasets suffer from a sim-to-real gap. Inspired by the Counterfactual World Model (CWM) paradigm, which can obtain point-wise correspondences by injecting a small tracer perturbation into a next-frame predictor and tracking its propagation, we extend this idea to generative video models for zero-shot flow extraction. We explore several popular architectures and find that successful zero-shot flow extraction in this manner is aided by three model properties: (1) distributional prediction of future frames (avoiding blurry or noisy outputs); (2) factorized latents that treat each spatio-temporal patch independently; and (3) random-access decoding that can condition on any subset of future pixels. These properties are uniquely present in the recently introduced Local Random Access Sequence (LRAS) architecture. Building on LRAS, we propose KL-tracing: a novel test-time inference procedure that injects a localized perturbation into the first frame, rolls out the model one step, and computes the Kullback–Leibler divergence between perturbed and unperturbed predictive distributions. Without any flow-specific fine-tuning, our method is competitive with state-of-the-art, task-specific models on the real-world TAP-Vid DAVIS benchmark and the synthetic TAP-Vid Kubric. Our results show that counterfactual prompting of controllable generative video models is an effective alternative to supervised or photometric-loss methods for high-quality flow.  [^1] Project website at: https://neuroailab.github.io/projects/kl_tracing/
@@
Figure 1: We introduce a zero-shot test-time inference procedure called KL-tracing, which extracts robust optical flow and point tracking from a generative video model on challenging in-the-wild videos. In every column, the green line links the query location in the first frame (top) to the position predicted by our method in the second frame (bottom). All clips are real-world internet videos and contain phenomena that classical, appearance-based optical flow methods find challenging: (A) Newton’s cradle, where both frames have four balls in the middle, but the balls are different; the example involves physical reasoning. (B) Globe has challenging in-place object rotation and the query point is in the textureless ocean. (C) Dog weaving through occluding poles with large, rapid motion, including depth changes and motion blur. (D) Soccer tackle with fast, diagonal motion with motion blur and partial occlusion. (E) Windmill rotation where the repetitive blades and uniform sky make local matching challenging. These examples highlight the benefits of leveraging a powerful video model to extract optical flow for challenging real-world scene dynamics.
@@ 1 Introduction
Extracting motion information (optical flow) from videos is a fundamental yet open challenge in computer vision with many applications. Due to the intractable cost of obtaining ground-truth labels from real-world videos, most supervised baselines [teed_raft_2020, wang_sea-raft_2024] are trained on synthetic datasets, e.g., FlyingChairs [fischer_flownet_2015], FlyingThings [mayer_large_2016] and Sintel [hutchison_naturalistic_2012]. While invaluable for progress, these datasets cover a narrow slice of motion statistics–predominantly rigid objects, limited lighting variation, and short temporal horizons. As a result, they under-represent the long-tail of real-world phenomena such as non-rigid deformation, atmospheric effects, rapid camera shake and textureless regions. Self-supervised models [jiang_doduo_2023, stone_smurf_2021] that can be trained on real videos attempt to bridge this gap, but they often rely on task-specific heuristics, such as photometric consistency or smoothness which fail under complex lighting, occlusion, or long-range motion. Consequently, both supervised and self-supervised optical flow baselines struggle to generalize to challenging in-the-wild videos.
@@
Inspired by successes across vision and language where large general-purpose models outperform smaller task-specific ones, we explore large-scale video models as a possible solution. Trained on massive repositories of real-world data, modern video models already demonstrate strong scene understanding [blattmann_stable_2023, nvidia_cosmos_2025, assran2025vjepa2selfsupervisedvideo, madan2024foundationmodelsvideounderstanding], suggesting an implicit grasp of optical flow [bear_unifying_2023, venkatesh_counterfactual_2023]. However, prior work extracting visual intermediates such as depth and illumination from these models still required supervised fine-tuning [shao2024chronodepth, fang2025relightvid]. Extending that recipe to optical flow would again depend on synthetic labels, facing the same sim-to-real domain gap. This motivates our search for a zero-shot procedure that can extract accurate optical flow from off-the-shelf video models without any additional training.
@@
In fact, such a procedure has been proposed in the Counterfactual World Model (CWM) framework [bear_unifying_2023]. Zero-shot optical flow is obtained by adding a small “tracer” perturbation to the source frame and tracking how a pretrained next-frame predictor propagates the tracer to the target frame (Section 2). In practice, however, because deterministic video models like CWM can only predict a single future state, it encourages predictions that average over future possibilities, yielding perceptually blurry frames that wash out the injected tracer, leading to less precise motion estimates (Section 4.1).
@@
To overcome this, we implement the perturb-and-track method with generative video models, which generate crisp predictions as they sample from a distribution instead of regressing to a mean. This is insufficient, however, as each state-of-the-art generative video model faces its own flow extraction challenges. Stable Video Diffusion (SVD) [blattmann_stable_2023] produces photorealistic frames, but its conditioning is weak and non-localized: generation is guided by a single global latent inverted from the target frame, so pixel-level edits introduce noisy differences, corrupting the extracted flow (Section 4.2). The recent Cosmos model [nvidia_cosmos_2025] can condition generation on ground-truth patches from the target frame, but its autoregressive rollout is strictly in raster order, so these provided patches must all lie in the raster top-left region of the image, far less informative than the same number of patches distributed randomly across the image, thus failing to accurately reconstruct the target frame (Section 4.3).
@@
Through a systematic analysis of state-of-the-art video models, we find three key properties for accurate zero-shot optical flow extraction: (1) distributional prediction of future frames (avoiding blurry or noisy outputs); (2) local tokenization, which encodes each spatio-temporal patch independently; and (3) random access decoding that can condition on any subset of future pixels. These are present in the recent Local Random Access Sequence (LRAS) architecture [lee_3d_2025], whose patch-level conditioning and random-order decoding provide much more fine-grained controllability than previous generative video models (Section 4.4). In addition, we propose a new method, KL-tracing, that further improves upon the probe-and-track method by using the autoregressive nature of LRAS which exposes the probability distributions for predictions. KL-tracing computes the perturbation difference in the prediction logit space, thereby bypassing noisy RGB differences resulting from sampling randomness. We achieve state-of-the-art results on the challenging real-world TAP-Vid DAVIS dataset and the synthetic TAP-Vid Kubric dataset relative to supervised and unsupervised flow baselines (Section 6).
@@ 2 Related Work
Optical Flow refers to estimating the per-pixel 2D motion between a pair of video frames. Most models [teed_raft_2020, wang_sea-raft_2024] are supervised on synthetic video datasets [fischer_flownet_2015, mayer_large_2016, hutchison_naturalistic_2012], posing a sim-to-real gap, e.g., predominantly rigid objects, limited lighting variation, and short temporal horizons, failing to capture the long tail of real-world phenomena such as non-rigid deformation, atmospheric effects, fast camera shakes, and fine-scale textureless regions. Alternatively, self-supervised models [jonschkowski_what_2020, liu_learning_2020, stone_smurf_2021] train on real-world videos, but often rely on task-specific architectures and heuristics such as photometric consistency and smoothness loss [jonschkowski_what_2020, liu_learning_2020, stone_smurf_2021]. Regardless, these optical flow methods are designed for short frame gaps and relatively non-complex scenes with little motion dynamics.
@@
Point Tracking follows a set of points across longer time horizons. Most methods are supervised [doersch_tap-vid_2023, harley_particle_2022, karaev_cotracker_2024], often on synthetic datasets, facing the sim-to-real gap. Alternatively, self-supervised methods rely on task-specific heuristics such as smoothness losses and cycle consistency [jabri_space-time_2020, bian_learning_2022, shrivastava_self-supervised_2024], i.e., tracking a point forward in time, then backward, should return to its original position. In contrast, we use a zero-shot method to extract flow from general-purpose self-supervised video models trained on diverse, large-scale datasets, as they better learn challenging real-world dynamics (Section 6). Further, the best performing trackers on long horizon benchmarks often exploit multi-frame context, which can obscure a model’s ability to resolve core frame-to-frame dynamics across variable timesteps. Therefore, we compare our zero-shot method primarily against baselines following the traditional, two-frame setting for flow.
@@
Deterministic Video Models predict a single future frame/latent  [hafner_learning_2019, hafner_dream_2020, hafner_mastering_2022, bear_unifying_2023, bardes_revisiting_2024]. Trained with either latent or pixel $\ell_{1}/\ell_{2}$ reconstruction losses, they are implicitly optimized to output the expectation over plausible futures. Whenever the conditioning signal—a single frame $F_{1}$, multiple frames $(F_{-N}\!\ldots\!F_{1})$, or a partially masked target $F_{2}^{\text{masked}}$—can lead to multiple future possibilities, the model averages them, producing spatially blurred predictions which hurt flow estimation.
@@
Generative Video Models. Generative models can sample different image or video generations from a probabilistic distribution. Diffusion models iteratively denoise a single noise sample into a generation [ho_denoising_2020, song_score-based_2021, saharia_photorealistic_2022, ho_video_2022, blattmann_stable_2023]. Autoregressive models [oord_conditional_2016, chen_pixelsnail_2017, razavi_generating_2019, clark_adversarial_2019] generate outputs sequentially, modeling the joint distribution over pixels or tokens, each element conditioned on previously generated elements. However, many of these models, e.g., Cosmos [nvidia_cosmos_2025], predict tokens in raster order and use a global tokenizer where each image patch is encoded with information from other patches. Hence, they lack fine-grained controllability which makes them challenging to use for flow extraction (Section 4.3). In contrast, a recent class of models, Local Random Access Sequence (LRAS), has several key properties which make them suitable for optical flow extraction (Section 4.4). Applying our novel test-time inference procedure to LRAS achieves state-of-the-art results (Section 6).
Tracks from Diffusion Models. Related to our work, there has been recent interest in using the emergent motion understanding capabilities of generative video models. The focus of these investigations has primarily been on reading out motion information from intermediate representations of diffusion models [tang2023emergentcorrespondenceimagediffusion, nam2025emergenttemporalcorrespondencesvideo]. In this work, we instead rely on a different extraction procedure (namely, propagating intervention points for tracking), which we find to be more transferable to generative architectures beyond diffusion. Our approach, when paired with design decisions that maximize controllability and locality, outperforms the latest diffusion-specific methods for flow (Section 6).
@@ 3 Test-time inference procedure and evaluation setup for optical flow
Figure 2: Test-time inference procedure for extracting flow from a pre-trained, frozen, generative video model, based on the Counterfactual World Model (CWM) paper [bear_unifying_2023]. This involves three steps: (1) Perturbation: add a small, white-colored 2D Gaussian dot perturbation to frame 1 at the location of the point we wish to track. (2) Generate model predictions conditioned on the two frames. For CWM, Cosmos, and LRAS, we provide frame 1 and masked patches of frame 2 (Sections 4.1, 4.3, 4.4). For Stable Video Diffusion, we provide the noised latents of both frames (Section 4.2). (3) Estimate optical flow by computing the RGB difference between the clean and perturbed predictions.
We use a zero-shot procedure to extract optical flow from various pre-trained video models. For all models, we utilize the generic method template illustrated in Figure 2, based on the Counterfactual World Model (CWM) paper [bear_unifying_2023]. Step 1. Inject a small perturbation. We duplicate the initial frame $F_{1}$ and perturb it with a small “white bump” to form $\tilde{F}_{1}$, i.e., a Gaussian centered at the query location $x_{q}$ with amplitudes 255 on each RGB channel and standard deviation $\sigma$ equal to 2.0. Step 2. Run model twice. Both clean and perturbed initial frames are separately forwarded through the frozen model, each time with a sparse mask of the second frame $({F}_{2}^{\text{masked}})$. Hence, two forward passes: $({F}_{1},{F}_{2}^{\text{masked}}\rightarrow{F}_{2}^{\text{pred}})$ and $(\tilde{F}_{1},{F}_{2}^{\text{masked}}\rightarrow\tilde{F}_{2}^{\text{pred}})$. Step 3. Estimate optical flow. Compute the RGB difference between the two predictions, $F_{2}^{\text{pred}}$ and $\tilde{F}_{2}^{\text{pred}}$. Then, take the arg-max to identify the location to which the perturbation was carried.
@@
Evaluation setup. We use TAP-Vid DAVIS [doersch_tap-vid_2023] and Kubric [greff_kubric_2022] for evaluation. TAP-Vid DAVIS contains real-world videos with human-annotated flow and occlusion labels, while Kubric is a synthetic dataset with ground-truth labels. Metrics. We use the following metrics from TAP-Vid [doersch_tap-vid_2023]: (1) Average Distance (AD) (or endpoint error (EPE)), the Euclidean distance between the predicted and ground-truth flow; (2) Average Jaccard (AJ), the “true positives” divided by “true positives” plus “false positive” plus “false negatives”, averaged over various thresholds; (3) $\delta^{x}_{\text{avg}}\!$, the fraction of points that are within a threshold of the ground truth location, averaged over various thresholds; and (4) Occlusion Accuracy (OA) for predicting occluded points.
@@ 4 Model evaluations
We conduct several studies extracting optical flow from various video model classes, highlighting issues we found with each. We investigate deterministic models (Section 4.1), diffusion models (Section 4.2), and autoregressive models (Section 4.3). Finally, we identify the key properties of generative models that aid precise flow prediction, and a model class that satisfies them (Section 4.4).
4.1 Study 1: Deterministic models produce blurry-averaged predictions
Method: Deterministic video models such as the Counterfactual World Model (CWM) [bear_unifying_2023, venkatesh_counterfactual_2023] predict a single future state. CWM is a visual foundation model that can be zero-shot prompted to perform many tasks, such as predicting flow, keypoints, object segments, counterfactuals, and depth. CWM was trained with an $\ell_{2}$ loss to perform masked next-frame prediction: $({F}_{1},F_{2}^{\text{masked}}\rightarrow{F}_{2}^{\text{pred}})$. It deterministically predicts a single future state ${F}_{2}^{\text{pred}}$, as it lacks a mechanism for sampling. To extract flow from CWM, we use the procedure in Section 2 with a small, red perturbation, following [bear_unifying_2023].